X^2+8x-8y+24=8 how do you identify the conic
Answer by lwsshak3(11628)

X^2+8x-8y+24=8 how do you identify the conic
complete the square
(x^2+8x+16)-8y=8+16-24+8y
(x+4)^2=8y
This is an equation of a parabola that opens upward with vertex at (-4,0)
Its basic form: (x-h)^2=4p(y-k), (h,k)=(x,y) coordinates of the vertex.